# Beamforming and Resource Allocation for Delay Optimization in RIS-Assisted OFDM Systems

### v3

@@ Abstract
This paper investigates a joint beamforming and resource allocation problem in downlink reconfigurable intelligent surface (RIS)-assisted orthogonal frequency division multiplexing (OFDM) systems to optimize average delay, where data packets for each user arrive at the base station stochastically. The sequential optimization problem is inherently a Markov decision process (MDP), thus falling within the scope of reinforcement learning. To effectively handle the mixed action space and reduce the state space dimensionality, a hybrid deep reinforcement learning (DRL) approach is proposed. Specifically, proximal policy optimization (PPO)-$\Theta$ is employed to optimize the RIS phase shift design, while PPO-$N$ is responsible for subcarrier allocation decisions. The active beamforming at the base station is then derived from the jointly optimized RIS phase shifts and subcarrier allocation decisions. To further mitigate the curse of dimensionality associated with subcarrier allocation, a multi-agent strategy is introduced to optimize subcarrier allocation indicators more efficiently. Moreover, to achieve more adaptive resource allocation and accurately capture network dynamics, key factors closely related to average delay—such as the number of backlogged packets in buffers and current packet arrivals—are incorporated into the state space. Furthermore, a transfer learning framework is introduced to enhance training efficiency and accelerate convergence. Simulation results demonstrate that the proposed algorithm significantly reduces average delay, enhances resource allocation efficiency, and achieves superior system robustness and fairness compared to baseline methods.
@@
Reconfigurable intelligent surface, deep reinforcement learning, OFDM, delay optimization.
@@ I Introduction
With the continuous evolution of wireless communication technologies, sixth-generation (6G) wireless communication systems are envisioned to achieve unprecedented performance in terms of capacity, latency, reliability, and energy efficiency[1]. To meet the stringent demands of future applications, extensive research is being conducted on enhancing the fundamental capabilities of wireless systems[2]. Among the various enabling technologies, reconfigurable intelligent surfaces (RIS) have emerged as a key enabler[3], drawing significant research interest due to their potential to reconfigure the wireless propagation environment and support enhanced system performance.
As a critical performance metric, latency plays a vital role in determining the overall performance of wireless systems, particularly for mission-critical applications such as remote surgery, autonomous driving, factory automation, and intelligent transportation systems [4]. These applications often require real-time data exchange. In such scenarios, even a small increase in latency can lead to severe consequences, such as system failures, road accidents, or disrupted services. Therefore, ensuring ultra-low latency, often on the order of tens of milliseconds or less [5], is essential for numerous mission-critical applications in the 6G era.
However, given the inherently limited nature of wireless resources, ensuring low-latency communication in RIS-assisted wireless systems remains a fundamental challenge, where the phase design and resource allocation are coupled. In addition to the challenge caused by joint optimization of phase design and resource allocation, limited spectrum and power resources constrain the system’s ability to support high data rates and real-time responsiveness, multi-path effects introduce rapid fluctuations in channel conditions and signal quality [6], which necessiates network management to be adaptive to the environment. On the other hand, different users may have different numbers of backlogged packets, which needs to be taken into account in delay-aware resource allocation [7]. Therefore, in delay-aware RIS-assisted wireless networks, it is critical to make sequential resource allocation and phase design decisions while being adaptive to real-time channel conditions and queue status [8], which is challenging due to complex network environment and mixed decision variables including continuous RIS phase shifts design and discrete subcarrier.
@@ I-A Prior Works and Motivation
The delay refers to the duration from the moment the packet is generated at the transmitter to its reception by the receiver. This metric is crucial for evaluating the performance of communication networks, especially for applications requiring real-time responsiveness. Numerous studies [9, 10, 11] have shown that queuing delay constitutes the dominant component of overall latency compared to other components, such as processing latency, transmission latency, and propagation latency.
@@
To address delay-aware resource allocation, some works leverage large-scale fading for spectrum and power management. For instance, in device-to-device (D2D)-enabled vehicular networks, resource allocation based on large-scale fading improves vehicle-to-infrastructure (V2I) ergodic capacity while ensuring vehicle-to-vehicle (V2V) reliability [12]. Similarly, single-input multiple-output (SIMO) systems have adopted power and rate allocation strategies using average signal and interference statistics under reliability and energy efficiency constraints [13]. Sun et al. further proposed a resource block and power control algorithm for D2D-based V2V communications without relying on instantaneous channel state information (CSI) [14]. However, these static methods struggle to adapt to rapid channel variations. To address this, Lyapunov optimization has gained traction for real-time delay-aware resource management, accounting for queue dynamics, channel fluctuations, and system constraints. For example, Guo et al. developed a cross-layer Lyapunov framework combining transport-layer rate control and physical-layer allocation to meet delay targets [15]. Similar approaches have led to robust scheduling policies that ensure queue stability and service quality in dynamic environments [16, 17, 18].
@@
Beyond conventional optimization techniques, deep reinforcement learning (DRL)-based algorithms have emerged as powerful tools for addressing delay-aware sequential resource allocation in wireless communication systems, offering a favorable balance between performance and computational complexity [19]. Leveraging the Markov decision process (MDP) framework, DRL is particularly well-suited for real-time, sequential resource allocation decisions. Early efforts in this domain, such as the Q-learning-based link scheduling algorithm proposed in [20], demonstrate how queue length, channel conditions, and energy constraints can be jointly considered to optimize average transmission delay. As the complexity of network environments increases, research has evolved from single-agent to multi-agent paradigms. For instance, multi-agent deep reinforcement learning (MADRL) has been adopted to better handle distributed decision-making scenarios. In [21], each V2V link is modeled as an autonomous agent that dynamically performs spectrum sharing and power control, effectively reducing latency while minimizing interference with V2I communications.
@@
While RIS has enhanced spectral efficiency, coverage, and energy efficiency [22, 23, 24, 25, 26], its application in latency-sensitive scenarios remains limited. Most existing works focus on mobile edge computing and neglect queuing delay, which is crucial for real-time communication [27, 28, 29]. For instance, [27] minimized task delay and energy in RIS-assisted non-orthogonal multiple access networks, and [28] addressed joint delay and rate in multiple-input multiple-output mobile edge computing systems. To address this gap, recent studies adopt Lyapunov-based methods for queue-aware resource allocation, converting long-term delay goals into per-slot decisions to adapt to dynamic traffic [30, 31]. For example, [31] jointly optimized power, channel assignment, and reflection coefficients in a semi-grant-free RIS-assisted system, improving queue stability over RIS-free baselines. However, current approaches remain insufficient. Many rely on complex alternating optimization and fail to explicitly include queuing delay in the objective, limiting their effectiveness in truly delay-sensitive RIS-assisted systems.
@@
This research aims to bridge this gap by proposing a novel approach that integrates RIS technology with advanced optimization methods, including DRL techniques, to optimize average delay and enhance network performance. Our approach provides a new perspective on addressing latency challenges in RIS-assisted systems and offers a fresh direction for improving system performance in delay-sensitive applications. This work seeks to unlock the untapped potential of RIS in managing queuing delay in practical communication systems, ultimately enhancing the efficiency of RIS-assisted wireless networks.
@@ I-B Contributions and Organization
This paper aims to investigate resource allocation methods for delay optimization in RIS-assisted networks by leveraging DRL, with a particular focus on maintaining user fairness. To achieve this goal, a hybrid DRL framework is proposed to tackle the challenge of delay optimization in such networks. The main contributions of this work are summarized as follows.
  • A hybrid DRL framework is adopted, where the base station (BS) controller acts as the agent and is trained using proximal policy optimization (PPO). Specifically, PPO-$\Theta$ with continuous action space is used for RIS phase shift optimization, while PPO-$N$ with discrete action space handles subcarrier allocation. To address the curse of dimensionality, a multi-agent strategy is employed, where each agent is responsible for the allocation of a single subcarrier.
@@
  • We evaluate the proposed method across various RIS-assisted communication network configurations, demonstrating its superiority over baseline methods. Simulation results indicate that the method significantly reduces average delay while ensuring user fairness, demonstrating its superiority over baseline methods. Specifically, fairness is enhanced in the balanced accumulation of data packets across user buffers, even under varying arrival rates. Moreover, the method exhibits excellent robustness, as it maintains stable performance under different delay taps, user distributions, and Rician fading factors.
The rest of the paper is organized as follows. In Section II, the system model is introduced, followed by the formulation of the joint RIS phase and resource allocation optimization problem. At the end of this section, the optimization problem is formally defined. In Section III elaborates our propoosed hybrid DRL-based scheme, including the overall framework, PPO-$\Theta$ design, PPO-$N$ design, beamforming design, and trainning methodology. Section IV presents numerical simulations and complexity analysis to validate the proposed method. Finally, conclusions are presented in Section V.
(a) Illustration of a RIS-assisted OFDM-MISO system communication scenario.
(b) Queueing model at the base station.
@@ II System Model and Problem formulation
We consider an RIS-assisted downlink multiple-input single-output (MISO) orthogonal frequency division multiplexing (OFDM) communication system, as illustrated in Fig. 1LABEL:sub@Fig:RIS, where a BS equipped with $N_{t}$ antennas serves $K$ single-antenna users. An RIS comprising $M$ passive reflecting elements is deployed to effectively enhance the transmission quality. The entire communication duration is divided into $T_{\mathrm{sum}}$ timeslots, indexed by $\{1,2,\ldots,t,\ldots,T_{\mathrm{sum}}\}$, each with a fixed length of $T$. Within each time slot, the wireless channel is assumed to remain approximately static. Each user’s data packets, which arrive randomly at the corresponding buffer at the BS, follow a first-come, first-served (FCFS) discipline, as illustrated in Fig. 1LABEL:sub@Fig:queue. In each time slot, the accumulated packets in the buffer $k$ are scheduled for transmission to the user $k$. Packets that can not be delivered will remain in the buffer and continue to accumulate, resulting in increased packet delay over time.
@@ II-A Signal Model
In the MISO-OFDM system considered in this paper, $\tilde{\mathbf{h}}^{(d)}_{l,k}[t]\in\mathbb{C}^{1\times N_{t}},l=0,1,...,L_{0}-1$ represents the time-domain baseband equivalent channel of the direct link from the BS to the user $k$ at the time slot $t$, where $L_{0}$ denotes the number of tap delays. Similarly, $\mathbf{\tilde{G}}_{l}[t]\in\mathbb{C}^{M\times N_{t}},l=0,1,...,L_{1}-1$ and $\tilde{\mathbf{r}}_{l,k}[t]\in\mathbb{C}^{1\times M},l=0,1,...,L_{2}-1$ respectively denote the time-domain baseband equivalent channels from the BS to the RIS and from the RIS to the user $k$ at the time slot $t$, where $L_{1}$ and $L_{2}$ are their respective numbers of tap delays. Therefore, the total maximum number of delay taps is $L=\operatorname*{max}\big{\{}L_{0},L_{1}+L_{2}-1\big{\}}$. Thus, the overall channel from the BS to the user $k$ at the time slot $t$ can be expressed as
@@
where $\tilde{\mathbf{G}}_{l}[t]=\mathbf{0}\text{,}\;l\in\begin{Bmatrix}1-L_{2}$\text{\,} $\ldots,-1\end{Bmatrix}\cup\begin{Bmatrix}L_{1},\ldots,L-1\end{Bmatrix}$, and $\boldsymbol{\Phi}[t]=\operatorname{diag}(e^{j\theta_{1}[t]},e^{j\theta_{2}[t]},...,e^{j\theta_{M}[t]})\in\mathbb{C}^{M\times M}$ is the RIS reflection phase shifts matrix at the time slot $t$. Let $\boldsymbol{\mathbf{\varphi}}[t]=[e^{j\theta_{1}[t]},e^{j\theta_{2}[t]},...,e^{j\theta_{M}[t]}]^{T}\in\mathbb{C}^{M\times 1}$ denote the RIS reflection phase shifts vector at the time slot $t$, where $\theta_{m}[t]$ refers to the phase shift of the reflection element $m$ at the time slot $t$. We set $\tilde{\mathbf{h}}^{(r)}_{l,k}[t]=\sum_{i=0}^{L_{2}-1}\mathrm{diag}(\tilde{\mathbf{r}}_{i,k}[t])\tilde{\mathbf{G}}_{l-i}[t]$, thus $\tilde{\mathbf{h}}_{l,k}[t]$ can further be expressed as
@@
where $\bar{\mathbf{h}}^{(d)}_{n,k}[t]$ and $\bar{\mathbf{h}}^{(r)}_{n,k}[t]$ represent the direct link channel and cascaded reflection channel of the user $k$ on subcarrier $n$ at the time slot $t$ in the frequency domain, respectively. Therefore the received signal of the user $k$ on subcarrier $n$ at the time slot $t$ can be represented as
$y_{n,k}[t]=(\bar{\mathbf{h}}^{(d)}_{n,k}[t]+\boldsymbol{\varphi}^{T}\bar{\mathbf{h}}^{(r)}_{n,k}[t])\mathbf{w}_{n}[t]s_{n}[t]+\nu_{n}[t],$ (4)
where $\mathbf{w}_{n}[t]\in\mathbb{C}^{N_{t}\times 1}$ denotes the beamforming vector on subcarrier $n$ at the time slot $t$, $s_{n}[t]$ represents the transmitted signal on subcarrier $n$ at the time slot $t$ with $\mathbb{E}\{|s_{n}[t]|^{2}\}=1$, and $\nu_{n}[t]$ is additive Gaussian white noise, which satisfies $\nu_{n}[t]\sim\mathcal{CN}(0,\sigma^{2})$. Thus, the signal-to-noise ratio (SNR) of the user $k$ on subcarrier $n$ at the time slot $t$ can be expressed as
@@
Therefore, the achievable rate for the user $k$ at the time slot $t$ is given by
@@
where $W$ represents the bandwidth of the subcarrier, and $\alpha_{n,k}[t]$ denotes the subcarrier allocation indicater at the time slot $t$. In particular, we have $\alpha_{n,k}[t]=1$ if subcarrier $n$ is allocated to the user $k$ at the time slot $t$, and $\alpha_{n,k}[t]=0$ otherwise.

### v4

# Beamforming and Resource Allocation for Delay Minimization in RIS-Assisted OFDM Systems

@@ Abstract
This paper investigates a joint beamforming and resource allocation problem in downlink reconfigurable intelligent surface (RIS)-assisted orthogonal frequency division multiplexing (OFDM) systems to minimize the average delay, where data packets for each user arrive at the base station (BS) stochastically. The sequential optimization problem is inherently a Markov decision process (MDP), thus falling within the remit of reinforcement learning. To effectively handle the mixed action space and reduce the state space dimensionality, a hybrid deep reinforcement learning (DRL) approach is proposed. Specifically, proximal policy optimization (PPO)-$\Theta$ is employed to optimize the RIS phase shift design, while PPO-$N$ is responsible for subcarrier allocation decisions. The active beamforming at the BS is then derived from the jointly optimized RIS phase shifts and subcarrier allocation decisions. To further mitigate the curse of dimensionality associated with subcarrier allocation, a multi-agent strategy is introduced to optimize the subcarrier allocation indicators more efficiently. Moreover, to achieve more adaptive resource allocation and accurately capture the network dynamics, key factors closely related to average delay—such as the number of backlogged packets in buffers and current packet arrivals are incorporated into the state space. Furthermore, a transfer learning framework is introduced to enhance the training efficiency and accelerate convergence. Simulation results demonstrate that the proposed algorithm significantly reduces the average delay, enhances the resource allocation efficiency, and achieves superior system robustness and fairness compared to baseline methods.
@@
Deep reinforcement learning, delay optimization, OFDM, reconfigurable intelligent surface.
@@ I Introduction
With the continuous evolution of wireless communication technologies, sixth-generation (6G) wireless communication systems are envisioned to achieve unprecedented performance in terms of capacity, latency, reliability, and energy efficiency[1]. To meet the stringent demands of future applications, extensive research is being conducted on enhancing the fundamental capabilities of wireless systems [2]. Among the various enabling technologies, reconfigurable intelligent surfaces (RISs) have emerged as a key enabler[3], drawing significant research interest due to their potential to reconfigure the wireless propagation environment [4] and support enhanced system performance [5, 6].
As a critical performance metric, latency plays a vital role in determining the overall performance of wireless systems, particularly for mission-critical applications, such as remote surgery, autonomous driving, factory automation, and intelligent transportation systems [7]. These applications often require real-time data exchange. In such scenarios, even a small increase in latency can lead to severe implications, such as system failures, road accidents, or disrupted services. Therefore, ensuring ultra-low latency, often on the order of tens of milliseconds or less [8], is essential for numerous mission-critical applications in the 6G.
However, given the inherently limited nature of wireless resources, ensuring low-latency communication in RIS-assisted wireless systems remains a fundamental challenge, where the phase design and resource allocation are coupled. In addition to the above mentioned challenge, limited spectrum and power resources constrain the system’s ability to support high data rates and real-time responsiveness, while multi-path effects introduce rapid fluctuations in the channel conditions and signal quality [9], which necessiates network management to be adaptive to the environment. On the other hand, different users may have different numbers of backlogged packets, which needs to be taken into account in delay-aware resource allocation [10]. Therefore, in delay-aware RIS-assisted wireless networks, it is critical to make sequential resource allocation and phase design decisions while being adaptive to real-time channel conditions and queue statuses [11], which is challenging due to complex network environments and mixed decision variables, including continuous RIS phase shift design and discrete variations of subcarrier allocation indicators.
@@ I-A Prior Works and Motivation
This section provides an overview of related work, with a summarized comparison shown in Table I. The delay refers to the duration from the instance the packet is generated at the transmitter to its reception by the receiver. This metric is crucial for evaluating the performance of communication networks, especially for applications requiring real-time responsiveness. Numerous studies (e.g., [12, 13, 14]) have shown that queuing delay constitutes the dominant component of the overall latency compared to other components, such as processing latency, transmission latency, and propagation latency.
@@
To address delay-aware resource allocation, some works leverage large-scale fading for spectrum and power management. For instance, in device-to-device (D2D)-enabled vehicular networks, resource allocation based on large-scale fading improves vehicle-to-infrastructure (V2I) ergodic capacity, while ensuring vehicle-to-vehicle (V2V) reliability [15]. Similarly, single-input multiple-output (SIMO) systems have adopted power and rate allocation strategies using average signal and interference statistics under reliability and energy efficiency constraints [16]. Sun et al. further proposed a resource block and power control algorithm for D2D-based V2V communications without relying on instantaneous channel state information (CSI) [17]. However, these static methods struggle to adapt to rapid channel variations. To address this, Lyapunov optimization has gained traction for real-time delay-aware resource management, accounting for queue dynamics, channel fluctuations, and system constraints. For example, Guo et al. developed a cross-layer Lyapunov framework combining transport-layer rate control and physical-layer allocation to meet delay targets [18]. Similar approaches have led to robust scheduling policies that ensure queue stability and service quality in dynamic environments [19, 20, 21].
TABLE I: Comparison of delay-oriented resource allocation methods
Research work RIS Delay-oriented RL Multi-agent Queuing delay
[18, 19, 20, 21]
[22]
[23]
[24]
[25]
The proposed work
@@
Beyond conventional optimization techniques, deep reinforcement learning (DRL)-based algorithms have emerged as powerful tools for addressing delay-aware sequential resource allocation in wireless communication systems, offering a favorable balance between performance and computational complexity [26]. Leveraging the Markov decision process (MDP) framework, DRL is particularly well-suited for real-time, sequential resource allocation decisions. Early efforts in this domain, such as the Q-learning-based link scheduling algorithm proposed in [22], demonstrate how queue length, channel conditions, and energy constraints can be jointly considered to minimize the average transmission delay. As the complexity of network environments increases, research has shifted from single-agent to multi-agent paradigms. For instance, multi-agent deep reinforcement learning (MADRL) has been adopted to better handle distributed decision-making scenarios. In [27], each V2V link was modeled as an autonomous agent that dynamically performs spectrum sharing and power control, effectively reducing latency while minimizing interference within V2I communications.
@@
While RISs can theoretically alter the spectral efficiency, coverage, and energy efficiency [28, 29, 30, 31, 32], their application in latency-sensitive scenarios remains limited. Most existing works focus on mobile edge computing and neglect queuing delay, which is crucial for real-time communication [23, 24, 33]. For instance, [23] minimized the task delay and energy in RIS-assisted non-orthogonal multiple access networks, while [24] addressed joint delay and rate in multiple-input multiple-output mobile edge computing systems. To address this gap, recent studies adopt Lyapunov-based methods for queue-aware resource allocation, converting long-term delay goals into per-slot decisions to adapt to dynamic traffic [34, 25]. For example, [25] jointly optimized the power, channel assignment, and reflection coefficients in a semi-grant-free RIS-assisted system, improving the queue stability over RIS-free baselines. However, current approaches remain insufficient. Many rely on complex alternating optimizations and fail to explicitly include queuing delay in their objective functions, limiting their effectiveness in truly delay-sensitive RIS-assisted systems.
@@
Our research aims to bridge this gap by proposing a novel approach that integrates the RIS technology with advanced optimization methods, including DRL techniques, to minimize the average delay and enhance the network performance. Our approach provides a new perspective on addressing latency challenges in RIS-assisted systems and offers a fresh direction for improving the system performance in delay-sensitive applications. In general, this work seeks to unlock the untapped potential of RISs in managing queuing delay in practical communication systems, ultimately enhancing the efficiency of RIS-assisted wireless networks.
@@ I-B Contributions and Organization
This paper aims to investigate resource allocation methods for delay minimization in RIS-assisted networks by leveraging DRL, with a particular focus on maintaining user fairness. To achieve this goal, a hybrid DRL framework is proposed to tackle the challenge of delay optimization in such networks. The main contributions of this work are summarized as follows.
  • A hybrid DRL framework is adopted, where the base station (BS) controller acts as the agent and is trained using proximal policy optimization (PPO). Specifically, PPO-$\Theta$ with continuous action space is used for the RIS phase shift optimization, while PPO-$N$ with discrete action space handles the subcarrier allocation. To address the curse of dimensionality, a multi-agent strategy is employed, where each agent is responsible for the allocation of a single subcarrier.
@@
  • We evaluate the proposed method across various RIS-assisted communication network configurations, demonstrating its superiority over baseline methods. Simulation results indicate that the method significantly reduces the average delay while ensuring user fairness, demonstrating its superiority over baseline methods. Specifically, fairness is enhanced in the balanced accumulation of data packets across user buffers, even under varying arrival rates. Moreover, the method exhibits excellent robustness, as it maintains stable performance under different delay taps, user distributions, and Ricean fading factors.
The rest of the paper is organized as follows: In Section II, the system model is introduced, followed by the formulation of the joint RIS phase and resource allocation optimization problem. At the end of this section, the optimization problem is formally defined. Section III elaborates on our propoosed hybrid DRL-based scheme, including the overall framework, PPO-$\Theta$ design, PPO-$N$ design, beamforming design, and trainning methodology. Section IV presents the numerical simulations and complexity analysis to validate the proposed method. Finally, conclusions are presented in Section V.
(a) Illustration of a RIS-assisted OFDM-MISO system communication scenario
(b) Queueing model at the BS
@@ II System Model and Problem formulation
We consider an RIS-assisted downlink multiple-input single-output (MISO) orthogonal frequency division multiplexing (OFDM) communication system, as illustrated in Fig. 1LABEL:sub@Fig:RIS, where a BS equipped with $N_{t}$ antennas serves $K$ single-antenna users. An RIS comprising $M$ passive reflecting elements is deployed to effectively enhance the transmission quality. The entire communication duration is divided into $T_{\mathrm{sum}}$ timeslots, indexed by $\{1,2,\ldots,t,\ldots,T_{\mathrm{sum}}\}$, each with a fixed length of $T$. Within each time slot, the wireless channel is assumed to remain approximately static. Each user’s data packets, which arrive randomly at the corresponding buffer at the BS, follow a first-come, first-served (FCFS) discipline, as illustrated in Fig. 1LABEL:sub@Fig:queue. In each time slot, the accumulated packets in buffer $k$ are scheduled for transmission to user $k$. Packets that cannot be delivered will remain in the buffer and continue to accumulate, resulting in increased packet delay over time.
@@ II-A Signal Model
In the MISO-OFDM system considered in this paper, $\tilde{\mathbf{h}}^{(d)}_{l,k}[t]\in\mathbb{C}^{1\times N_{t}},l=0,1,...,L_{0}-1$ represents the time-domain baseband equivalent channel of the direct link from the BS to user $k$ at time slot $t$, where $L_{0}$ denotes the number of tap delays. Similarly, $\mathbf{\tilde{G}}_{l}[t]\in\mathbb{C}^{M\times N_{t}},l=0,1,...,L_{1}-1$ and $\tilde{\mathbf{r}}_{l,k}[t]\in\mathbb{C}^{1\times M},l=0,1,...,L_{2}-1$ respectively denote the time-domain baseband equivalent channels from the BS to the RIS and from the RIS to user $k$ at time slot $t$, where $L_{1}$ and $L_{2}$ are their respective numbers of delay taps. Therefore, the total maximum number of delay taps is $L=\operatorname*{max}\big{\{}L_{0},L_{1}+L_{2}-1\big{\}}$. Thus, the overall channel from the BS to user $k$ at time slot $t$ can be expressed as
@@
where $\tilde{\mathbf{G}}_{l}[t]=\mathbf{0}\text{,}\;l\in\begin{Bmatrix}1-L_{2}$\text{\,} $\ldots,-1\end{Bmatrix}\cup\begin{Bmatrix}L_{1},\ldots,L-1\end{Bmatrix}$, and $\boldsymbol{\Phi}[t]=\operatorname{diag}(e^{j\theta_{1}[t]},e^{j\theta_{2}[t]},...,e^{j\theta_{M}[t]})\in\mathbb{C}^{M\times M}$ is the RIS reflection phase shift matrix at time slot $t$. Let $\boldsymbol{\mathbf{\varphi}}[t]=[e^{j\theta_{1}[t]},e^{j\theta_{2}[t]},...,e^{j\theta_{M}[t]}]^{T}\in\mathbb{C}^{M\times 1}$ denote the RIS reflection phase shift vector at time slot $t$, where $\theta_{m}[t]$ refers to the phase shift of the reflection element $m$ at time slot $t$. We set $\tilde{\mathbf{H}}^{(r)}_{l,k}[t]=\sum_{i=0}^{L_{2}-1}\mathrm{diag}(\tilde{\mathbf{r}}_{i,k}[t])\tilde{\mathbf{G}}_{l-i}[t]$; thus, $\tilde{\mathbf{h}}_{l,k}[t]$ can be further expressed as
@@
where $\bar{\mathbf{h}}^{(d)}_{n,k}[t]$ and $\bar{\mathbf{H}}^{(r)}_{n,k}[t]$ represent the direct link channel and cascaded reflection channel of user $k$ on subcarrier $n$ at time slot $t$ in the frequency domain, respectively. Therefore, the received signal of user $k$ on subcarrier $n$ at time slot $t$ can be represented as
$y_{n,k}[t]=\big{(}\bar{\mathbf{h}}^{(d)}_{n,k}[t]+\boldsymbol{\varphi}^{T}\bar{\mathbf{H}}^{(r)}_{n,k}[t]\big{)}\mathbf{w}_{n}[t]s_{n}[t]+\nu_{n}[t],$ (4)
where $\mathbf{w}_{n}[t]\in\mathbb{C}^{N_{t}\times 1}$ denotes the beamforming vector on subcarrier $n$ at time slot $t$, $s_{n}[t]$ represents the transmitted signal on subcarrier $n$ at time slot $t$ with $\mathbb{E}\{|s_{n}[t]|^{2}\}=1$, while $\nu_{n}[t]$ is additive Gaussian white noise, which satisfies $\nu_{n}[t]\sim\mathcal{CN}(0,\sigma^{2})$. Thus, the signal-to-noise ratio (SNR) of user $k$ on subcarrier $n$ at time slot $t$ can be expressed as
@@
Therefore, the achievable rate for user $k$ at time slot $t$ is given by
@@
where $W$ represents the bandwidth of the subcarrier, and $\alpha_{n,k}[t]$ denotes the subcarrier allocation indicator at time slot $t$. In particular, we have $\alpha_{n,k}[t]=1$ if subcarrier $n$ is allocated to user $k$ at time slot $t$, and $\alpha_{n,k}[t]=0$ otherwise.